OpenStudy (anonymous):

Rigorously prove that a finite sum of rationals cannot evaluate to an irrational.

OpenStudy (agreene):

Oh set theory... I used to know this proof off the back of my hand, gimme a min to remember it--or forget it, you never know.

OpenStudy (agreene):

actually, this is not true.

OpenStudy (anonymous):

Alright, then, the question becomes rigorously prove that this isn't true. XD I guess by contrapositive assume.

OpenStudy (agreene):

i need to remember how you do that whole, series adding to different sums thing, lol

OpenStudy (agreene):

there is some sum I learned a while back that naturally sums to e, but if you add anything to it it becomes any other number you want, bleh

OpenStudy (anonymous):

Well, because we have Euler's relationship, and we have the infinite sum of 1/n^2 converging at pi something, it shouldn't be too hard to evaluate the famous irrationals. No, the issue is we need to prove, for all cases, that either a finite sum of rationals can evaluate to an irrational, or not.

OpenStudy (agreene):

oh, wait you said a finite sum, that was an infinite sum... i revert back to saying it's true

OpenStudy (anonymous):

Yeah, I can't think of anything but a limit creating an irrational.

OpenStudy (agreene):

well, since all rational numbers can be expressed as a fraction a/b... and irrationals cannot be expressed as a/b thats probably where the proof lies.

OpenStudy (agreene):

a,b are integers.

OpenStudy (agreene):

Let x,y ∈ ℚ. Then x = a/b and y = c/d for a,b,c,d ∈ ℤ. Then x + y = (a/b) + (c/d) = (ad + bc) / (bd) = p/q for some p,q ∈ ℤ. Therefore (x + y) ∈ ℚ, i.e. the rationals are closed under addition.

OpenStudy (anonymous):

Makes sense. I suppose I was hoping for a roundabout proof so that it could help me with a previous question. XD The question was, "Rigorously demonstrate the conditions that must be satisfied such that the limit of an infinite sum of rationals converge to an irrational." Which I had no idea how to start, so I decided to start by proving everything around it.

OpenStudy (agreene):

these are some really annoying questions. is this for like an algebraic structures or a proofs class?

OpenStudy (anonymous):

Something like that. Informal challenges given out by the department once in a while.

OpenStudy (anonymous):

Don't worry; I'm not cheating. XD

OpenStudy (agreene):

even worse.

OpenStudy (agreene):

this is bringing back bad memories of weeks spent in the library doing my undergrad research :P

OpenStudy (agreene):

so, as i mentioned, rationals must be able to satisfy a/b in order to become irrational this must no longer be able to happen. I'm sure the proof lies in an annoying limit dealing with this.

OpenStudy (anonymous):

Bleaargghh. My school department is so helpful, giving out problems like this, but not the solutions.